OpenStudy (anonymous):

how do you solve x^2-2x-33=2

OpenStudy (anonymous):

I think you may have to use quadratic equation

OpenStudy (anonymous):

no nee for the quadratic equation here, x^2-2x-33=2; -2 x^2-2x-35=0 (x+5)(x-7)=0

OpenStudy (dumbcow):

Bring 2 over by subtracting 2 on both sides x^2-2x-35 =0 Look for factors of -35 that add up to -2 As shown above :)